# HRI-RECAPP-2024-05

### v1

@@ IV.1 The Overlapping Solution
In case of overlapping solution, the signal strength corresponding to the $\tau^{+}\tau^{-}$ and $\gamma\gamma$ channels receive substantial contribution from both the CP-even and CP-odd states simultaneously [^2] Here we have considered CP-conserving potential. As a result, the $h$ and $A$ states would not interfere with each other.. In contrast, for the $b\bar{b}$ mode only the CP-even state contribute as the trilinear coupling $AZZ$ is zero at tree level. As a result, the signal strength can be expressed in the following manner -
@@
$m_{H}$          $125$
$m_{h}$           $92-98$
$m_{A}$          $92-98$
$m_{H^{\pm}}$          $80-230$
$\tan{\beta}$          $0.5-100$
$\sin({\beta-\alpha})$          $-0.66-0.62$
@@
Table 2: The scan range which is used for the MC sampling. All the Masses are in GeV.
@@
We generated MC samples in the scan range as described in Table  2. After testing them against various theoretical and experimental constraints, the allowed parameter space is illustrated in Figure 1. The region of the parameter space that pass the theoretical constraints are depicted by the blue points while the region that pass the experimental constraints are depicted by red points. The plot illustrates that a nearly degenerate solution with both the CP-even and CP-odd Higgs in the $92-98$ GeV mass range is viable for the charged Higgs mass ranging around $195160<m_{H}^{\pm}<$ 195. We will use the overlapping region of the two coloured point distributions to test the aforementioned anomalies.
@@
Figure 2: Correlations amongst the signal strengths in Eq. (6). Here, the total $\chi^{2}$ is displayed using the colour bar and the best fit point is given by star marker. The ATLAS and CMS results with their corresponding $1\sigma$ band is also represented.
@@
Figure  2 illustrates the total chi-square distribution for points that are compatible with all three anomalies and the best fit point is marked by a star, corresponding to $\chi^{2}_{\rm min}$. The experimentally observed signal strength with their $1\sigma$ band is also superimposed in the plot to test them against this model’s best fit point. The chi-square fit reveals two distinct branches, differentiated by the sign of $\sin(\beta-\alpha)$. The main branch, characterized by positive values, is densely populated, while the second branch, which has negative values, is sparsely populated due to the elimination of many points by various constraints. The figure  3 also illustrates the sign of the electroweak coupling parameter. The best fit point indicated by a red star has a large positive $\sin(\beta-\alpha)$ value, while the vector boson coupling with the SM Higgs boson, indicated by the color bar, is weakened in that case.
@@
Given that the di-photon excess is most pronounced around $95$ GeV, we also plot the allowed points within the $94-96$ GeV range over the CMS and ATLAS results for the signal strength in the ${\gamma\gamma}$ channel, as shown in Figure  4. The expected and observed CMS limits are shown by the black dashed and solid lines. The green and yellow bands indicate the $1\sigma$ and $2\sigma$ uncertainties and the plot is overlaid by the ATLAS observed $95\%$ confidence level limits on the signal strength with the red dashed and solid lines. The combined signal strength (CMS+ATLAS) at $95.4$ GeV with its error bar is also shown using a red dot. The points explaining the anomaly at the $1\sigma$ level, for $3$ degrees of freedom, corresponding to the $\gamma\gamma$, $\tau\tau$ and $b\bar{b}$ channels as shown in equation  9, which requires $\chi^{2}<3.53$, are plotted in dark red while the points explaining it at the $3\sigma$ level, demanding $\chi^{2}<7.8147$ are shown in peru color (Less likely points are given in sky blue). The best fit point in the $94-96$ range is also indicated using midnight blue color. It can be clearly seen that the parameter points are suited to explain the excesses observed. The details of the best fit points as marked in the two figures are shown in Table 3.
Figure 3: $\sin(\beta-\alpha)$ plotted against $\chi^{2}_{\rm sum}$. The $\chi^{2}_{\rm min}$ is indicated by a star.
@@
Fig 2       125.0       92.23       92.21       176.81       3.26       0.36       1.1255       0.2395       0.1164       0.02235
Fig 4       125.0       94.03       95.71       191.77       3.22       0.38       1.107       0.2510       0.1228       0.0598
Table 3: BPs extracted from Figures 2 and 4. Masses are in GeV.
@@ IV.2 The Single Solution
$m_{H}$          $125$
$m_{h}$           $92-97$
$m_{A}$          $80-230$
$m_{H^{\pm}}$          $80-230$
$\tan{\beta}$          $0.5-100$
$\sin({\beta-\alpha})$          $-0.66-0.62$
Table 4: The scan range which is used for the MC sampling. All the Masses are in GeV.
In this case, only the $h$ state is responsible for explaining all the three anomalies. We sampled points for the scan range described in Table  4, and the points that pass the various constraints are depicted in Figure 5, wherein the blue shaded region indicates the points that pass the various theoretical constraints while the red shaded region indicates region of parameter space that survives after imposing different experimental bounds. The plots clearly represent the allowed parameter space for the CP-odd and Charged Higgs masses, given that we fix our CP-even mass to lie in the range $92-97$ GeV. Though the charged Higgs mass is bounded at around $160$ GeV, the CP-odd pseudoscalar covers almost the entire scan range. Note that the allowed CP-odd scalar mass also lie in the $90-100$ GeV mass window, hinting to the overlapping solution that we discussed in the previous section. We move ahead with testing the overlapping points with the observed anomalies.
@@
The total $\chi^{2}$ fit for the points passing all the constraints is displayed in Figure 6, wherein the best fit point (i.e., again, the $\chi^{2}_{\rm min}$ one) is indicated with a star. We have again overlaid the plot with the experimentally observed data with the $1\sigma$ band. The plot clearly depicts that the best fit point lies within the experimental $1\sigma$ boundary. The two bands here again are differentiated with the sign of $\sin(\beta-\alpha)$, wherein the densely populated region is positive while the sparsely populated one negative. This is also depicted in Figure 8, where the best fit point is depicted by a red star, which has a large $\sin(\beta-\alpha)$ value, while the vector boson coupling with the SM Higgs boson is weakened, which is indicated by the color bar. The mass region lying between $94-96$ GeV is also displayed on top of the CMS and ATLAS results for the signal strength in the $\gamma\gamma$ channel in Figure 8 (with the same color coding as that in Figure 4). The figure clearly depicts the best fit point in that particular region to be close to the mean experimentally observed value. Finally the best fit BPs in these cases are shown in Table 5.
       $m_{H}$       $m_{h}$       $m_{A}$       $m_{H^{\pm}}$       $\tan\beta$       $\sin(\beta-\alpha)$       $\mu_{\tau^{+}\tau^{-}}(h)$       $\mu_{\gamma\gamma}(h)$       $\mu_{b\bar{b}}(h)$       $\chi^{2}_{\tau^{+}\tau^{-}+\gamma\gamma+b\bar{b}}$
Figure 6       125.0       92.52       124.73       181.77       3.11       0.41       0.8429       0.1942       0.1457       1.022
Figure 8       125.0       95.84       85.03       201.35       3.11       0.41       0.6631       0.2240       0.2351       1.122
Table 5: The details of the best fit points extracted from Figures 6 and 8. Masses are in GeV.
Figure 6: Correlations amongst the signal strengths in Eq. (6). Here, the total $\chi^{2}$ is displayed using the colour bar and the best fit point is given by star marker. ATLAS and CMS results with their corresponding $1\sigma$ band is also added.
Figure 7: $\sin(\beta-\alpha)$ plotted against $\chi^{2}_{\rm sum}$. The best fit point is indicated by a red star.
@@ V Conclusions
In summary, we have proven that somewhat anomalous data produced at LEP and the LHC in $b\bar{b}$ and $\tau^{+}\tau^{-}$ as well as $\gamma\gamma$ final states, respectively, all clustering in a 10 GeV or so mass window around 95 GeV, are consistent with the possibility of the 2HDM-Type I in inverted mass hierarchy explaining these. Specifically, two configurations are possible: one where both the (degenerate) $h$ and $A$ states cooperate to explain the aforementioned anomalies and another where only the $h$ states does so. This is an intriguing result, as such Higgs states can well be probed in collateral signatures specific to the 2HDM Type-I in inverted mass hierarchy, as emphasised in various previous literature [63, 64, 65, 66, 67, 68, 69, 70]. To aid testing this theoretical hypothesis, we have produced two pairs of BPs amenable to further phenomenological analysis, each pair corresponding to one of the above two solutions, wherein the parameter space point giving the best fit to all anomalies.
@@ Acknowledgments
The work of S.M. is supported in part through the NExT Institute and the STFC Consolidated Grant No. ST/ L000296/1. A.S acknowledges the support from SERB-National Postdoctoral fellowship (Ref No: PDF/2023/002572). A.K. acknowledges the support from Director’s Fellowship at IIT Gandhinagar. All authors thank Tanmoy Mondal and Prasenjit Sanyal for their help in discovering a mistake in their calculations.

### v2

@@ IV.1 The Overlapping Solution
In the case of the overlapping solution, the signal strengths corresponding to the $\tau^{+}\tau^{-}$ and $\gamma\gamma$ channels receive substantial contribution from both the CP-even and CP-odd states simultaneously [^2] Here we have considered a CP-conserving potential. As a result, the $h$ and $A$ states would not interfere with each other.. In contrast, for the $b\bar{b}$ mode only the CP-even state contribute as the trilinear coupling $AZZ$ is zero at tree level. As a result, the signal strengths can be expressed in the following manner:
@@
$m_{H}$    $125$ GeV
$m_{h}$    $94~\text{GeV}-96~\text{GeV}$
$m_{A}$    $94~\text{GeV}-96~\text{GeV}$
$m_{H^{\pm}}$    $140~\text{GeV}-250~\text{GeV}$
$\tan{\beta}$    0.51000.5-100
$\sin({\beta-\alpha})$    0.600.60-0.60-0.60
@@
Table 2: The scan range which is used for the MC sampling for the overlapping solution.
Figure 1: Results of the scan in Table 2 mapped against the Higgs boson masses. The blue regions are allowed by stability, unitarity and perturbativity constraints while the red regions are allowed by Higgs data, $b\rightarrow s\gamma$ and EW precision constraints.
@@
We generated MC samples in the scan range as described in Table  2. After testing them against various theoretical and experimental constraints, the allowed parameter space is illustrated in Figure 1. The region of the parameter space that pass the theoretical constraints are depicted by the blue points while the region that pass the experimental constraints are depicted by red points. The plot illustrates that a nearly degenerate solution with both the CP-even and CP-odd Higgs in the 949694-96 GeV mass range is viable for the charged Higgs mass ranging over the interval $152~\text{GeV}<m_{H}^{\pm}<168~\text{GeV}$. We will use the overlapping region of the two coloured point distributions to test the aforementioned anomalies.
@@
Figure 2: Correlations amongst the signal strengths in Eq. (6). Here, the total $\chi^{2}$ is displayed using the colour bar and the best fit point is given by the star marker. The ATLAS and CMS results with their corresponding $1\sigma$ band are also represented.
@@
Figure 2 illustrates the total $\chi^{2}$ distribution for points that are compatible with all three anomalies and the best fit point is marked by a star, corresponding to $\chi^{2}_{\rm min}$. The experimentally observed signal strength with their $1\sigma$ band is also superimposed in the plot to test them against this model’s best fit point. The chi-square fit reveals two distinct branches, differentiated by the sign of $\sin(\beta-\alpha)$. The longer branch is characterised by the positive values, while the smaller one by the negative values. Figure 3 further highlights the sign of the EW coupling parameter, with the best-fit point in both branches again marked by a red star. Notably, the positive branch contains a best fit point that simultaneously explains all the three anomalies within the $1\sigma$ region. However, an analysis by the ATLAS collaboration on the 2HDM Type-I [66] imposes significant restrictions on the positive region of the value $\sin(\beta-\alpha)$, limiting its viability. The negative branch also allows for a simultaneous explanation of all the three anomalies within the $1\sigma$ region. However, when the $b\bar{b}$ channel anomaly is examined in isolation, it cannot be explained at the $1\sigma$ level. While the vector boson coupling with the SM Higgs boson, indicated by the colour bar, is weakened as in this case, it remains within the allowed range.
@@
Given that the di-photon excess is most pronounced around $95$ GeV, we also plot the allowed points within the 949694-96 GeV range over the CMS and ATLAS results for the signal strength in the $\gamma\gamma$} channel, for the negative branch, as shown in Figure  4. The expected and observed CMS limits are shown by the black dashed and solid lines. The green and yellow bands indicate the $1\sigma$ and $2\sigma$ uncertainties and the plot is overlaid by the ATLAS observed $95\%$ C.L. limits on the signal strength with the red dashed and solid lines. The combined signal strength (CMS+ATLAS) at $95.4$ GeV with its error bar is also shown using a red dot. The points explaining the anomaly at the $1\sigma$ level, for $3$ degrees of freedom, corresponding to the $\gamma\gamma$, $\tau\tau$ and $b\bar{b}$ channels as shown in Eq. (9), which requires $\chi^{2}<3.53$, are plotted in dark red while the points explaining it at the $2\sigma$ level, demanding $\chi^{2}<7.8147$, are shown in peru colour. (Less likely points are given in sky blue.) The best fit point in the 949694-96 range is also indicated using midnight blue color. It can be clearly seen that many parameter points are suited to explain the excesses observed. The details of the best fit point as marked in the figure is shown in Table 3. Although we have imposed a global $\chi^{2}$ goodness of test at $95\%$ C.L., the best-fit BP listed in Table  3 yields a $\chi^{2}$ value of $171.14$, , which also falls within the $3\sigma$ region corresponding to the more stringent HiggsSignals constraint, $\chi^{2}\leq\chi^{2}_{\rm SM}+11.83$.
Figure 3: The value of $\sin(\beta-\alpha)$ plotted against $\chi^{2}_{\rm sum}$. The $\chi^{2}_{\rm min}$ is indicated by a star. The color gradient represents the strength of the coupling between the SM like Higgs boson and the pair of vector bosons.
@@
Figure 2    125.0    94.36    94.44    166.63    2.33    0.17-0.17    0.836    0.296    0.027    3.402
Table 3: BP extracted from Figure 2 corresponding to the negative region of $\sin(\beta-\alpha)$. The masses of the Higgs states are given in units of GeV. The positive region of $\sin(\beta-\alpha)$ is strongly constrained by the ATLAS analysis of the 2HDM Type-I  [66], rendering it less viable.
@@ IV.2 The Single Solution
$m_{H}$    $125$ GeV
$m_{h}$    $94~\text{GeV}-96~\text{GeV}$
$m_{A}$    $90~\text{GeV}-250~\text{GeV}$
$m_{H^{\pm}}$    $140~\text{GeV}-250~\text{GeV}$
$\tan{\beta}$    0.51000.5-100
$\sin({\beta-\alpha})$    0.600.60-0.60-0.60
Table 4: The scan range which is used for the MC sampling for the single solution.
In this case, only the $h$ state is responsible for explaining all the three anomalies. We sampled points for the scan ranges described in Table  4 and the points that pass the various constraints are depicted in Figure 5, wherein the blue shaded region indicates the points that pass the various theoretical constraints while the red shaded region indicates the part of parameter space that survives after imposing different experimental bounds. The plots clearly represent the allowed parameter space for the CP-odd and charged Higgs masses, given that we fix our CP-even mass to lie in the range 949694-96 GeV. Although the charged Higgs mass is bounded at around $155$ GeV, the CP-odd one covers almost the entire scan range. The plots also clearly illustrate the degeneracy between the CP-odd and the charged Higgs states, which is required to satisfy the EW precision constraints. Note that the allowed CP-odd scalar mass also lies within the $94~\text{GeV}~\text{to}~96$ GeV mass window, hinting to the overlapping solution that we discussed in the previous section. We move ahead with testing the allowed points with the observed anomalies.
@@
The total $\chi^{2}$ fit for the points passing all the constraints is displayed in Figure 6, wherein the best fit point (i.e., again, the $\chi^{2}_{\rm min}$ one) is indicated with a star. We have again overlaid the plot with the experimentally observed data within the $1\sigma$ band. Here again the branches are differentiated by the sign of $\sin(\beta-\alpha)$, with the larger branch depicting the positive while the smaller one depicting the negative values. The plot demonstrates that, while a simultaneous solution exists for all three cases in the positive branch, no point in the negative branch satisfies all three anomalies simultaneously within the $2\sigma$ level ($\chi^{2}<7.8147$), once the ATLAS constraint is imposed. Nevertheless, viable points do appear in the negative branch when the requirement is relaxed to the $3\sigma$ level ($\chi^{2}<14.156$). Figure 7 provides additional insight into the sign of the electroweak coupling parameter, with red stars denoting the best-fit points in each of the two branches.
Figure 6: Correlations amongst the signal strengths in Eq. (6). Here, the total $\chi^{2}$ is displayed using the colour bar and the best fit point is given by star marker. ATLAS and CMS results with their corresponding $1\sigma$ bands are also added.
Figure 7: The value of $\sin(\beta-\alpha)$ plotted against $\chi^{2}_{\rm sum}$. The best fit point is indicated by a star. The color gradient represents the strength of the coupling between the SM like 125 GeV Higgs boson and the pair of vector bosons.
@@ V Conclusions
In summary, we have proven that somewhat anomalous data produced at LEP and the LHC in $b\bar{b}$ as well as $\tau^{+}\tau^{-}$ and $\gamma\gamma$ final states, respectively, all clustering in a mass window around 95 GeV, are consistent with the possibility of the 2HDM Type-I in inverted mass hierarchy (i.e., $m_{H}=125$ GeV) explaining these through one of its neutral Higgs states. Specifically, only one configuration is possible: the one where both the (degenerate) $h$ and $A$ states cooperate to explain the aforementioned anomalies, if we take the ATLAS result into account. This is an intriguing finding, as such a Higgs state can also be probed in collateral signatures specific to the 2HDM Type-I in inverted mass hierarchy, as emphasised in various previous literature [67, 68, 69, 70, 71, 72, 73, 74]. To aid testing this theoretical hypothesis, we have produced a BP amenable to further phenomenological analysis, wherein the parameter space point is giving the best fit to all anomalies. We also found that there does not exist a solution for only the CP-even case, that could explain the three anomalies simultaneously in the 2HDM Type-I even within $2\sigma$.
@@ Acknowledgments
The work of S.M. is supported in part through the NExT Institute and the STFC Consolidated Grant ST/X000583/1. A.S. acknowledges the support from SERB-National Postdoctoral fellowship (Ref No: PDF/2023/002572). A.K. acknowledges the support from Director’s Fellowship at IIT Gandhinagar. All authors thank Tanmoy Mondal and Prasenjit Sanyal for their help in discovering a mistake in their calculations.